State if it is Binomial, Hypergeometric, Geometric, Negative Binomial or Poisson: Five cards are drawn at random from a deck of cards for a poker hand. Find the probability that in that hand you have at least one diamond card.
Answer #1

this is hypergeometric distribution as events are not independent:

probability that in that hand you have at least one diamond card =1-P(none of the dimaond card)=1-39C5/52C5

=1-575757/2598960 =2023203/2598960 =0.778466